A train of mass 220 tonnes starts from rest up a grade of 1 in 90 incline. It reaches a speed of 72 km/hr in a distance of 900 m (up the incline). The acceleration is uniform and the tractive resistance is 0,75 % of the weight of the train. Use Newton’s second law to calculate the tractive effort required.


During your calculations clearly show:


a=?

TE=?

1
Expert's answer
2021-09-26T11:56:17-0400
"20^2=2(900)a\\a=0.22\frac{m}{s^2}\\"

"TE=220000(0.22+9.8(0.011)+9.8(0.0075))\\=88.3\ kN"